# Dingyi Zhao1, 7, Keming Wu6, Haozhe Wang8, Ping Nie9, Yan Teng1, 🖂, Yingchun Wang1

### v1

@@ Abstract
Artificial intelligence is undergoing the paradigm shift from closed language models to interconnected agent systems capable of external perception and information integration. As a representative embodiment, Deep Research Agents (DRAs) systematically exhibit the capabilities for task decomposition, cross-source retrieval, multi-stage reasoning, and structured output, which markedly enhance performance on complex and open-ended tasks. However, existing benchmarks remain deficient in evaluation dimensions, response formatting, and scoring mechanisms, limiting their capacity to assess such systems effectively. This paper introduces a rigorous benchmark and a multidimensional evaluation framework tailored to DRAs and report-style responses. The benchmark comprises 214 expert-curated challenging queries distributed across 10 broad thematic domains, each accompanied by manually constructed reference bundles to support composite evaluation. The framework enables comprehensive evaluation of long-form reports generated by DRAs, incorporating integrated scoring metrics for semantic quality, topical focus, and retrieval trustworthiness. Extensive experimentation confirms the superior performance of mainstream DRAs over web-search-tool-augmented reasoning models, yet reveals considerable scope for further improvement. This study provides a robust foundation for capability assessment, architectural refinement, and paradigm advancement in DRA systems.
@@ 1 Introduction
The core architecture of artificial intelligence paradigms is undergoing a transition from closed, static, parameter-driven Large Language Models (LLMs) to active, cognitive, and interconnected agent systems endowed with external perception and integrative mechanisms (Wang et al., 2024; Zhang et al., 2025b). Amid growing demands for heterogeneous information acquisition and increasing task complexity, Deep Research has gradually emerged as a paradigm of agent systems characterized by cognitive planning and integrative capabilities (Xu & Peng, 2025), which enable autonomous task decomposition, cross-source retrieval, multi-step reasoning, and structured expression, thereby substantially enhancing model adaptability and expressive performance in real-world applications (Huang et al., 2025; Zhang et al., 2025a).
@@
To address the limitations inherent in existing benchmarks and evaluation systems, our study proposes a rigorous benchmark, authored by human experts and specifically designed for DRAs, which targets high-difficulty and high-precision tasks with the goal of systematically assessing the overall performance in report-style long-text generation. On this basis, we develop a multidimensional evaluation framework that is intended to measure both the quality and the credibility of generated report-style responses. Our research yields three key contributions:
@@
    We introduce a rigorous benchmark paired with manually constructed reference bundles comprising query-specific and general-report rubrics, trustworthy source links, focus-anchor and focus-deviation keywords. Covering 214 challenging report-style entries across multiple domains, Rigorous Bench enables broad and granular characterization of DRA tasks, which addresses existing deficiencies in evaluation dimensions and response formats.
@@
    We develop a systematic and multidimensional evaluation framework for report-style outputs, which captures key processes of DRAs such as task reasoning, information retrieval, content synthesis, and structured articulation. By jointly modeling semantic quality, topical focus, and retrieval trustworthiness, our framework overcomes limitations of conventional methods and demonstrates high transferability to long-text generations beyond DRAs.
@@
    We conduct large-scale experiments involving five mainstream DRAs, one advanced agent model, and seven reasoning models enhanced with web-search tools. Quantitative results indicate that DRAs consistently outperform tool-augmented models in overall task execution and report generation quality, while revealing persistent limitations in architectural paradigms and behavioral mechanisms that warrant further refinement.
@@ 2.1 Deep Research Agents
Early LLMs, such as GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2023), rely on static training corpora and closed parameter spaces, while their knowledge is entirely derived from the training phase and cannot be updated or supplemented during inference. To overcome the epistemic constraints of static language models, researchers have explored mechanisms for integrating LLMs with external tools, giving rise to the paradigm of Tool-Augmented LLMs, exemplified by GPT-4 (Achiam et al., 2023), Gemini 1.5 (Google, 2024), Claude 3 (Anthropic, 2024), Toolformer (Schick et al., 2023), and Qwen 1.5 Agent (Alibaba, 2024). These models leverage external interfaces such as web browsers and code environments to enable dynamic information acquisition and cross-modal perception, reflecting a shift from knowledge encapsulation toward tool-mediated cognition.
@@ 2.2 Existing Benchmarks
With the growing adoption of Tool-Augmented LLMs, both academia and industry have increasingly focused on their performance in web-search tasks. Existing benchmarks, including GAIA (Mialon et al., 2023), WebWalker (Wu et al., 2025), BrowseComp (Wei et al., 2025), WideSearch (Wong et al., 2025), BrowseComp-Plus (Chen et al., 2025b), and DeepResearch Bench (Bosse et al., 2025), primarily evaluate LLMs using closed-form queries that require verifiable short answers to facilitate automated scoring and alignment. However, they lack coverage of key behaviors such as task decomposition, cross-source retrieval, and structured synthesis, and provide limited assessment of content hierarchy, discourse structure, and information integration. In addition, most rely on surface-level matching or similarity metrics, such as Exact Match, BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and BERTScore (Zhang et al., 2019), which struggle to capture semantic depth and structural fidelity, thereby limiting their effectiveness in evaluating the true capabilities of DRAs.
Contemporaneous work has begun to explore evaluation methods tailored to report-style outputs. DeepResearch Bench (Du et al., 2025) is the first to assess reference answers alignment and retrieval. However, it depends heavily on static reference reports, making it difficult to accommodate evolving query expectations. Its automated rubrics lack contextual sensitivity and focus on generic surface, failing to reflect human preferences for report quality and structure. Its retrieval evaluation emphasizes consistency between statements and cited links but overlooks the credibility of sources. Benchmarks such as ResearchQA (Yifei et al., 2025), DeepResearch Arena (Wan et al., 2025), and ReportBench (Li et al., 2025) constructed 21K, 10K, and 0.6K academic tasks respectively. These benchmarks also rely on automatically generated rubrics that exhibit limited stability and interpretability, which undermines their reliability for high-precision judgment. Additionally, large-scale benchmarks increase evaluation costs, reducing their practical utility for evaluating DRAs.
@@ 3.1 Domains
Figure 1: Distribution of benchmark entries.
The Rigorous Bench dataset was meticulously constructed and repeatedly validated by human experts, comprising 214 high-complexity entries across diverse domains. It is designed to challenge existing DRAs in task understanding, decomposition, execution, and aggregation. Based on semantic relevance, entries are systematically categorized into ten broad domains, with detailed distribution provided in Figure 1 and Appendix A.1. The dataset exhibits a high degree of professionalism and diversity in its design, effectively simulating complex and dynamic queries encountered in real-world scenarios, and demonstrating broad coverage and strong representativeness.
@@ 3.2.1 Query
To enhance consistency and reproducibility, Rigorous Bench systematically incorporates spatiotemporal robustness into the design of queries and rubrics. For fact-based tasks, each query explicitly defines temporal and geographic boundaries to mitigate external fluctuations in response content. For open-ended tasks, rubrics emphasize logical structure and reasoning validity to ensure generalizability and adaptability. All queries adopt the instruction “write a report/essay/…” to unify output format and reinforce task orientation. Overall, the query design achieves a high degree of semantic complexity, task diversity, and evaluative operability.
@@ 3.2.2 Query-Specific Rubrics
QSRs are designed to evaluate the task completion quality. Each QSR is custom-built by experts based on corresponding query, reflecting human expectations and evaluative preferences regarding factual accuracy and logical validity. Scoring follows a clearly defined binary (Yes/No) or ternary (Yes/Partial/No) scheme to ensure consistency and operational clarity. Each query is equipped with at least 8 QSRs, with a total score of 30, assigned in accordance with task importance.
QSRs are deeply embedded within the semantic structure of each task, offering high alignment and diagnostic precision. Their design spans core dimensions such as information coverage, mechanism explanation, structured expression, semantic precision, source verification, evidence organization, heterogeneity analysis, methodological transparency, temporal logic, and interdisciplinary integration. The QSRs provide both theoretical grounding and practical guidance for evaluating DRAs performance, while also laying a structural foundation for future automated scoring systems.
@@ 3.2.3 General-Report Rubrics
GRRs are designed to assess the quality of structured expression. Independent of specific queries, GRRs evaluate reports from a general perspective using binary judgment across seven key dimensions, namely structural organization, logical clarity & expression, informational coverage & content depth, citation quality & source credibility, originality & insight, data usage & analytical rigor, and formatting consistency. Comprising 48 rubrics with a total score of 73, the design emphasizes generalizability, normative clarity, and expressive strength, providing a unified quality standard for performance evaluation across tasks and models, as outlined in Appendix A.2.
@@ 3.3 Construction Pipeline
To ensure the difficulty, quality, and semantic validity, Rigorous Bench adopts a multi-stage review pipeline for systematic verification and refinement of expert-generated entries. This process integrates manual design, machine auditing, and cross-validation to establish a highly granular framework for data generation and quality control, as illustrated in the upper half of Figure 2.
In the initial stage, experts design diverse data units based on the unified Construction Guide and their own expertise within predefined domains. The units then undergo LLMs auditing to detect semantic inconsistencies, logical errors, and factual inaccuracies. During the first round of manual review, QSRs are verified for validity, with rubric criteria refined. For binary items with intermediate cases, a clearly defined “Partial” option is introduced to enhance scoring granularity and consistency. After preliminary filtering, queries and QSRs are standardized and rewritten to ensure stylistic and structural correctness. TSLs and Keywords are also supplemented and revised at this stage. The revised entries then undergo LLM-based difficulty testing, followed by the second round of manual review, which focuses on evaluating QSRs from observers’ perspective and assessing the anchoring effectiveness of keywords. Each link is individually verified for accessibility and authority. The third round of cross-review conducts a comprehensive quality check, covering domain classification, formatting standards, and structural integrity. Examples of entries can be found in Appendix A.3.
@@ 4 Evaluation Framework
Built on semantic quality, topical focus, and retrieval trustworthiness, a structured and multidimensional evaluation framework is proposed for report-style generation tasks, featuring an integrated scoring system inspired by principles from statistics and operations. It prioritizes transparency, interpretability, and practical utility for robust evaluation of complex tasks, as illustrated in Figure 2.
@@ 4.1 Semantic Quality
Semantic quality evaluates the overall performance of response reports in terms of task completion and general quality. It integrates scores from both QSRs and GRRs. Drawing on the Weighted Average Method (WAM) from Multi-Attribute Decision Making (MADM), the metric applies ratio normalization to both scores and assigns weighting coefficients $\alpha$ and $\beta$, where $\alpha+\beta=1$, to construct a fused model of multidimensional quality signals. The calculation formula is as follows:
$\text{Quality}=\alpha\cdot\mathcal{N}_{\text{Ratio}}\left[\ \sum_{i=1}^{N}\text{QSR}_{\text{score}}^{(i)}\ \right]+\beta\cdot\mathcal{N}_{\text{Ratio}}\left[\ \sum_{j=1}^{M}\text{GRR}_{\text{score}}^{(j)}\ \right]\in[0,1]$
where $\text{QSR}_{\text{score}}^{(i)}$ denotes the $i$-th QSR score; $\text{GRR}_{\text{score}}^{(j)}$ refers to the $j$-th GRR score; $N$ and $M$ represent the number of QSRs and GRRs respectively; $\mathcal{N}_{\text{Ratio}}[\cdot]$ denotes the ratio normalization; $\alpha$ and $\beta$ are weighting parameters reflecting the relative importance in the overall quality assessment.
@@ 4.2 Topical Focus
Topical focus is evaluated through the SemanticDrift metric, which jointly considers the absence of FAKs and the misuse of FDKs to measure the degree of thematic deviation in response reports.
$\text{FAK}_{\text{Drift}}$ quantifies the omission of core keywords. The frequency of each keyword in the report is scaled by its expected value $\epsilon$, and a minimum function is introduced to implement a threshold control. A FAK is penalized when its frequency falls below $\epsilon$, thereby enforcing the requirement for substantive keyword presence and ensuring that the score reflects semantic completeness rather than accidental occurrence. Drawing on the TF × IDF paradigm from information retrieval, $\text{FAK}_{\text{Drift}}$ is constructed as the product of frequency component and semantic relevance, which is defined as:
$\text{FAK}_{\text{Drift}}=1-\frac{1}{K}\sum_{k=1}^{K}\left[\min\left(\frac{\text{freq}^{(k)}}{\epsilon_{+}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(k)})\right]\in[0,1]$
where $\text{freq}^{(k)}$ is frequency of the $k$-th FAK; $\text{rele}^{(k)}\in[1,2,3,4,5]$ is the semantic relevance assessed by LLMs; $K$ represents the number of FAKs; $\epsilon_{+}$ is the expectation scaling factor. A higher $\text{FAK}_{\text{Drift}}$ value indicates weaker thematic focus and insufficient coverage of core concepts in the report. Similarly, $\text{FDK}_{\text{Drift}}$, which quantifies the intensity of thematic distraction, is defined as follows:
$\text{FDK}_{\text{Drift}}=\frac{1}{L}\sum_{l=1}^{L}\left[\min\left(\frac{\text{freq}^{(l)}}{\epsilon_{-}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(l)})\right]\in[0,1]$
where $\text{freq}^{(l)}$ is frequency of the $l$-th FDK; $\text{rele}^{(l)}\in[1,2,3,4,5]$ is the relevance score; $L$ represents the number of FDKs; $\epsilon_{-}$ is the scaling factor. A higher $\text{FDK}_{\text{Drift}}$ indicates stronger thematic deviation. The SemanticDrift is computed as a weighted combination of $\text{FAK}_{\text{Drift}}$ and $\text{FDK}_{\text{Drift}}$:

### v2

# from Answers to Reports

@@ Abstract
As an embodiment of intelligence evolution toward interconnected architectures, Deep Research Agents (DRAs) systematically exhibit the capabilities in task decomposition, cross-source retrieval, multi-stage reasoning, information integration, and structured output, which markedly enhance performance on complex and open-ended tasks. However, existing benchmarks remain deficient in evaluation dimensions, response format, and scoring mechanisms, limiting their effectiveness in assessing such agents. This paper introduces Dr. Bench, a multidimensional evaluation framework tailored to DRAs and long-form report-style responses. The benchmark comprises 214 expert-curated challenging tasks across 10 broad domains, each accompanied by manually constructed reference bundles to support composite evaluation. This framework incorporates metrics for semantic quality, topical focus, and retrieval trustworthiness, enabling a comprehensive evaluation of long reports generated by DRAs. Extensive experimentation confirms the superior performance of mainstream DRAs over web-search-tool-augmented reasoning models, yet reveals considerable scope for further improvement. This study provides a robust foundation for capability assessment, architectural refinement, and paradigm advancement of DRAs.
Yang Yao1, 2Yixu Wang1, 3Yuxuan Zhang4Yi Lu5Tianle Gu1, 6Lingyu Li1, 7
Dingyi Zhao1, 7Keming Wu6Haozhe Wang8Ping Nie9Yan Teng1, 🖂Yingchun Wang1
1 Shanghai Artificial Intelligence Laboratory 2 The University of Hong Kong 3 Fudan University
4 University of British Columbia 5 University of Toronto 6 Tsinghua University 7 Shanghai Jiao Tong University
8 Hong Kong University of Science and Technology 9 Peking University
yaoyangacademia@outlook.com, tengyan@pjlab.org.cn
@@ 1 Introduction
The core architecture of AI paradigms is undergoing a transition from closed, static, parameter-driven Large Language Models (LLMs) to active, cognitive, and interconnected agent systems endowed with external perception and integrative mechanisms (Wang et al., 2024; Zhang et al., 2025b). Amid growing demands for heterogeneous information acquisition and increasing task complexity, Deep Research has gradually emerged as a paradigm of agent systems characterized by cognitive planning and integrative capabilities (Xu and Peng, 2025), which enable autonomous task decomposition, cross-source retrieval, multi-step reasoning, and structured expression, thereby substantially enhancing model adaptability and expressive performance in real-world applications (Huang et al., 2025; Zhang et al., 2025a).
@@
To address the limitations inherent in existing evaluations, our study proposes a rigorous benchmark, authored by human experts and specifically designed for DRAs, which targets high-difficulty and high-precision tasks with the goal of systematically assessing the overall performance in report-style long-text generation. On this basis, we develop a multidimensional evaluation framework intending to accurately and stably measure the quality and credibility of generated reports. Our research yields three key contributions:
@@
    We introduce the D(eep)r(esearch). Bench paired with manually constructed reference bundles comprising query-specific and general-report rubrics, trustworthy sources, focus-anchor and focus-deviation keywords. Covering 214 challenging report-oriented entries across 10 broad domains, Dr. Bench rigorously and efficiently enables broad and granular characterization of DRA tasks, which addresses existing deficiencies in evaluation dimensions and response formats.
@@
    We develop a systematic and multidimensional evaluation framework for long-form report-style outputs, which captures key processes of DRAs such as task reasoning, information retrieval, content synthesis, and structured articulation. By jointly modeling semantic quality, topical focus, and retrieval trustworthiness, our framework overcomes limitations of conventional methods and demonstrates high transferability to long-text generations beyond DRAs.
@@
    We conduct large-scale experiments involving five mainstream DRAs, one advanced agent model, and seven reasoning models enhanced with web-search tools. Quantitative results indicate that DRAs consistently outperform tool-augmented models in overall task execution proficiency and report generation quality, while revealing persistent limitations in architectural paradigms and behavioral mechanisms that warrant further refinement and continued development.
@@ 2.1 Deep Research Agents
Early LLMs, such as GPT-3 (Brown et al., 2020) and PaLM (Chowdhery et al., 2023), rely on static training corpora and closed parameter spaces, while their knowledge is entirely derived from the training phase and cannot be updated or supplemented during inference. To overcome the epistemic constraints of static language models, researchers have explored mechanisms for integrating LLMs with external tools, giving rise to the paradigm of Tool-Augmented LLMs, exemplified by GPT-4 (Achiam et al., 2023), Gemini 1.5 (Google, 2024), Claude 3 (Anthropic, 2024), Toolformer (Schick et al., 2023), and Qwen 1.5 Agent (Alibaba, 2024). These models leverage external interfaces such as web browsers to enable dynamic information acquisition and cross-modal perception, reflecting a shift from knowledge encapsulation toward tool-mediated cognition.
@@ 2.2 Existing Benchmarks
With the growing adoption of Tool-Augmented LLMs, both academia and industry have increasingly focused on their performance in web-search tasks. Existing benchmarks, including GAIA (Mialon et al., 2023), WebWalker (Wu et al., 2025), BrowseComp (Wei et al., 2025), WideSearch (Wong et al., 2025), BrowseComp-Plus (Chen et al., 2025b), and Deep Research Bench (Bosse et al., 2025), primarily evaluate LLMs using closed-form queries that require verifiable short answers to facilitate automated scoring and alignment. However, they lack coverage of key behaviors such as task decomposition, cross-source retrieval, and structured synthesis, and provide limited assessment of content hierarchy, discourse structure, and information integration. A rough comparison of existing benchmarks can be found in Appendix F. In addition, most rely on surface-level matching or similarity metrics, such as Exact Match, BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and BERTScore (Zhang et al., 2019), which struggle to capture semantic depth and structural fidelity, thereby limiting their effectiveness in evaluating the true capabilities of DRAs.
Contemporaneous work has begun to explore evaluation methods tailored to report-style outputs. DeepResearch Bench (Du et al., 2025) is the first to assess reference answers alignment and retrieval. However, it depends heavily on static reference reports, making it difficult to accommodate evolving query expectations. Its automated rubrics lack contextual sensitivity and focus on generic surface, failing to reflect human preferences for report quality and structure. Its retrieval evaluation emphasizes consistency between statements and cited links but overlooks the credibility of sources. Benchmarks such as ResearchQA (Yifei et al., 2025), DeepResearch Arena (Wan et al., 2025), and ReportBench (Li et al., 2025) constructed 21K, 10K, and 0.6K academic tasks respectively. These benchmarks also rely on automatically generated rubrics that exhibit limited stability and interpretability, which undermines their reliability for high-precision judgment. Additionally, large-scale benchmarks increase evaluation costs, especially for expensive DRAs, reducing their practical utility.
@@ 3.1 Domains
The Dr. Bench was meticulously constructed and repeatedly validated by human experts, comprising 214 high-complexity entries across diverse domains. It is designed to challenge existing DRAs in task understanding, decomposition, execution, and aggregation. Based on semantic relevance, entries are systematically categorized into ten broad domains, with detailed distribution provided in Figure 1 and Appendix A. The dataset exhibits a high degree of professionalism and diversity in its design, effectively simulating complex and dynamic queries encountered in real-world scenarios, and demonstrating broad coverage and strong representativeness.
Figure 1: The proportion of Dr. Bench entries across domains.
@@ 3.2.1 Query
To enhance consistency and reproducibility, Dr. Bench systematically incorporates spatiotemporal robustness into the design of queries and rubrics. For fact-based tasks, each query explicitly defines temporal and geographic boundaries to mitigate external fluctuations in response content. For open-ended tasks, rubrics emphasize logical structure and reasoning validity to ensure generalizability and adaptability. All queries adopt the instruction “write a report/essay/…” to unify output format and reinforce task orientation. Overall, the query design achieves a high degree of semantic complexity, task diversity, and evaluative operability.
@@ 3.2.2 Query-Specific Rubrics
QSRs are designed to evaluate the task completion quality. Each QSR is custom-built by experts based on corresponding query, reflecting human expectations and evaluative preferences regarding factual accuracy and logical validity. Scoring follows a clearly defined binary (Yes/No) or ternary (Partial) scheme to ensure consistency and operational clarity. Each query is equipped with at least 8 QSRs, with a total score of 30, assigned in accordance with task importance.
QSRs are deeply embedded within the semantic structure of each task, offering high alignment and diagnostic precision. Their design spans core dimensions such as information coverage, mechanism explanation, structured expression, semantic precision, source verification, evidence organization, heterogeneity analysis, methodological transparency, temporal logic, and interdisciplinary integration. Appendix G outlines the core dimensions of QSRs design together with their corresponding descriptions. The QSRs provide both theoretical grounding and practical guidance for evaluating DRAs performance, while also laying a structural foundation for future automated scoring systems.
@@ 3.2.3 General-Report Rubrics
GRRs assess the quality of structured expression. Independent of specific queries, GRRs evaluate reports from a general perspective using binary judgment across 7 key dimensions, namely structural organization, logical clarity & expression, informational coverage & content depth, citation quality & source credibility, originality & insight, data usage & analytical rigor, and formatting consistency.
GRRs constitute a necessary, objective, and universally applicable set of standards that delineate the fundamental human expectations for report-style responses: a qualified report must not only accomplish the assigned task but also satisfy general quality requirements such as structural completeness, logical coherence, and insightful argumentation. Comprising 48 rubrics with a total score of 73, the design emphasizes generalizability, normative clarity, and expressive strength, thereby providing a unified quality benchmark for performance evaluation across tasks and models, and establishing a comprehensive evaluation framework grounded in the dual pillars of content and expression, as outlined in Appendix B.
@@ 3.2.4 Trustworthy-Source Links
To prevent TSLs expiration, we (1) defined stability and longevity requirements for queries and TSLs during data construction, mandating that TSLs originate from authoritative, official, and primary sources (e.g., government or academic institutions) that are reliably accessible and resistant to change, thereby ensuring stability from the outset; (2) introduced hostname matching at the metric mechanism to reduce dependence on specific pages, and incorporated hyperparameters in TrustworthyBoost, enabling full-match stability to be adjusted across evaluation targets; and (3) established long-term maintenance through regular link validity checks (e.g., quarterly) and the planned use of snapshots as permanent references.
@@ 3.2.6 Focus-Deviation Keywords
For example, when a query focuses on “men’s football matches”, the term “women” would divert the report off-topic. Similarly, when the query concerns the “electric vehicles”, references to “fuel cars” or “motorcycles” are classified as FDK. During data construction, these terms were annotated by experts and validated through multiple reviews to ensure their consistency and appropriateness.
@@ 3.3 Construction Pipeline
To ensure the difficulty, quality, and semantic validity, Dr. Bench adopts a multi-stage review pipeline for systematic verification and refinement of expert-generated entries. This process integrates manual design, machine auditing, and cross-validation to establish a highly granular framework for data generation and quality control, as illustrated in the upper half of Figure 2.
In the initial stage, experts design diverse data units based on the unified Construction Guide and their own expertise within predefined domains. The units then undergo LLMs auditing to detect semantic inconsistencies, logical errors, and factual inaccuracies. During the 1st round of manual review, QSRs are verified for validity, with rubric criteria refined. For binary items with intermediate cases, a clearly defined “Partial” option is introduced to enhance scoring granularity and consistency. After preliminary filtering, queries and QSRs are standardized and rewritten to ensure stylistic and structural correctness. TSLs and Keywords are also supplemented and revised at this stage. The revised entries then undergo LLM-based difficulty testing, followed by the 2nd round of manual review, which focuses on evaluating QSRs from observers’ perspective and assessing the anchoring effectiveness of keywords. Each link is individually verified for accessibility and authority. The 3rd round of cross-review conducts a comprehensive quality check, covering domain classification, formatting standards, and structural integrity. Examples can be found in Appendix C.
@@ 4 Evaluation Framework
Built on semantic quality, topical focus, and retrieval trustworthiness, a structured and multidimensional evaluation framework is proposed for report-style generation tasks, featuring an integrated scoring system inspired by principles from statistics and operations. It prioritizes transparency, interpretability, and practical utility for robust evaluation of complex tasks, as illustrated in the lower half of Figure 2.
@@ 4.1 Semantic Quality
Semantic quality evaluates the overall performance of response reports in terms of task completion and general quality. It integrates scores from both QSRs and GRRs. Drawing on the Weighted Average Method (WAM) from Multi-Attribute Decision Making (MADM), the metric applies ratio normalization to both scores and assigns weighting coefficients $\alpha$ and $\beta$, where $\alpha+\beta=1$, to construct a fused model of multidimensional quality signals. The calculation formula of Quality $\in[0,1]$ is as follows:
Quality $=\alpha\cdot\mathcal{N}_{\text{Ratio}}\left[\sum_{i=1}^{N}\text{QSR}_{\text{score}}^{(i)}\right]+\beta\cdot\mathcal{N}_{\text{Ratio}}\left[\sum_{j=1}^{M}\text{GRR}_{\text{score}}^{(j)}\right]$
where $\text{QSR}_{\text{score}}^{(i)}$ denotes the $i$-th QSR score; $\text{GRR}_{\text{score}}^{(j)}$ refers to the $j$-th GRR score; $N$ and $M$ represent the number of QSRs and GRRs respectively; $\mathcal{N}_{\text{Ratio}}[\cdot]=\cdot\,/\,\sum_{i}\mathrm{FullScore}(r_{i})$ denotes the ratio normalization; $\alpha$ and $\beta$ are weighting parameters reflecting the relative importance in the overall quality assessment.
@@ 4.2 Topical Focus
Topical focus is evaluated through the SemanticDrift metric, which jointly considers the absence of FAKs and the misuse of FDKs to measure the degree of thematic deviation.
$\text{FAK}_{\text{Drift}}\in[0,1]$ quantifies the omission of core keywords. The frequency of each keyword in the report is scaled by its expected value $\epsilon$, and a minimum function is introduced to implement a threshold control. A FAK is penalized when its frequency falls below $\epsilon$, thereby enforcing the requirement for substantive keyword presence and ensuring that the score reflects semantic completeness rather than accidental occurrence. Drawing on the TF × IDF paradigm from information retrieval, $\text{FAK}_{\text{Drift}}$ is constructed as the product of frequency component and semantic relevance:
$\text{FAK}_{\text{Drift}}=1-\frac{1}{K}\sum_{k=1}^{K}\left[\min\left(\frac{\text{freq}^{(k)}}{\epsilon_{+}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(k)})\right]$
where $\text{freq}^{(k)}$ is frequency of the $k$-th FAK; $\text{rele}^{(k)}\in\{1,2,3,4,5\}$ is the semantic relevance assessed by LLMs; $K$ represents the number of FAKs; $\epsilon_{+}$ is the expectation scaling factor. A higher $\text{FAK}_{\text{Drift}}$ value indicates weaker thematic focus and insufficient coverage of core concepts in the report. Similarly, $\text{FDK}_{\text{Drift}}\in[0,1]$, which quantifies the intensity of thematic distraction, is defined as follows:
$\text{FDK}_{\text{Drift}}=\frac{1}{L}\sum_{l=1}^{L}\left[\min\left(\frac{\text{freq}^{(l)}}{\epsilon_{-}},1\right)\times\mathcal{N}_{\text{Ratio}}(\text{rele}^{(l)})\right]$
where $\text{freq}^{(l)}$ is frequency of the $l$-th FDK; $\text{rele}^{(l)}\in\{1,2,3,4,5\}$ is the relevance score; $L$ represents the number of FDKs; $\epsilon_{-}$ is the scaling factor. A higher $\text{FDK}_{\text{Drift}}$ indicates stronger thematic deviation. The SemanticDrift is a weighted combination of $\text{FAK}_{\text{Drift}}$ and $\text{FDK}_{\text{Drift}}$: